# BRIGHT: A globally distributed multimodal building damage assessment dataset with very-high-resolution for all-weather disaster response

### v1

@@ 3.3 Model training
To train these models, we use a combination of cross-entropy loss and Lovasz softmax loss (Berman et al., 2018). Cross-entropy loss serves as the basic loss function for dense prediction tasks, while Lovasz softmax loss effectively addresses sample imbalance, between non-building and building pixels, and across different damage levels.
@@ 3.4 Accuracy assessment
We adopt overall accuracy (OA), F1 score (F1), and mean intersection over union (mIoU) to evaluate the performance of DL models on building damage assessment. These metrics are commonly used in previous research on building damage assessment (Zheng et al., 2021). Following the setup in prior unimodal building damage assessment studies and the related xView2 Challenge (Gupta et al., 2019), F1 score is used to assess the performance of DL models in the building localization and damage classification subtasks. Finally, OA and mIoU are employed to measure the overall quality of the final damage proxy map, providing a comprehensive assessment of the models’ ability to localize buildings and classify damage levels accurately.
4 Results and Discussions
4.1 BRIGHT evaluation results
Method | () | () | Final OA () | Final mIoU () | IoU per class () | |||
Background | Intact | Damaged | Destroyed | |||||
UNet | 89.19 | 54.69 | 94.73 | 59.74 | 96.16 | 69.84 | 17.64 | 55.32 |
DeepLabV3+ | 87.69 | 59.77 | 95.59 | 62.92 | 95.80 | 74.21 | 22.63 | 59.03 |
SiamAttnUNet | 88.90 | 58.93 | 95.47 | 62.76 | 96.13 | 73.94 | 19.91 | 61.08 |
SiamCRNN | 90.39 | 63.56 | 96.20 | 65.23 | 96.55 | 77.42 | 26.26 | 60.68 |
ChangeOS | 90.55 | 65.99 | 96.20 | 65.69 | 96.54 | 77.58 | 28.36 | 60.27 |
DamageFormer | 90.98 | 66.44 | 96.49 | 66.86 | 96.84 | 79.08 | 28.64 | 62.90 |
ChangeMamba | 91.60 | 67.93 | 96.65 | 67.19 | 97.07 | 80.16 | 30.35 | 61.19 |
Table 4 shows the evaluation results for each model on the test set. We observe that ChangeMamba achieves the best overall performance, with an OA of 96.65, an mIoU of 67.19, and the highest and scores (91.60 and 67.93, respectively). DamageFormer also performs well, closely following ChangeMamba, with an mIoU of 66.86 and an OA of 96.49. Both models demonstrate strong capability in both building localization and damage classification tasks. The accuracy advantage of ChangeMamba and DamageFormer underscores the importance of leveraging advanced deep learning architectures to improve performance in complex tasks such as building damage assessment. For models using the direct prediction approach (UNet, DeepLabV3+, SiamAttnUNet), DeepLabV3+ delivers the best results, with an mIoU of 62.92 and an OA of 95.59. However, its performance still lags behind that of the decoupled models, further emphasizing the advantage of task decoupling.
In order to ensure that the evaluation is not dominated by a few events with a large number of images (like Turkey-Earthquake-2023), Table 5 further presents the event-level mIoU for each model across 9 of 12 disaster events. ChangeMamba and DamageFormer achieve the highest average mIoU, with scores of 55.45 and 56.18, respectively. Notably, DamageFormer performs particularly well on events such as Goma-VE-2022, La Palma-VE-2021, and Derna-FL-2023, demonstrating its robustness across different types of disasters. While performance varies across events, earthquake-related events such as Les Cayes-EQ-2021, Morocco-EQ-2023, and Turkey-EQ-2023 present greater challenges for all models, where the average mIoU remains relatively low. This highlights the need for further research to improve model robustness for earthquake damage assessment, where damage patterns are often more complex and diverse.
@@ 4.1 BRIGHT evaluation results
Beirut-EP-2020 | 52.43 | 39.84 | 46.48 | 49.97 | 48.15 | 55.04 | 50.05 |
Bata-EP-2021 | 46.42 | 58.02 | 41.99 | 58.34 | 58.00 | 60.80 | 59.84 |
Goma-VE-2022 | 58.80 | 60.10 | 62.06 | 60.34 | 61.34 | 62.75 | 58.90 |
Les Cayes-EQ-2021 | 40.07 | 39.37 | 40.03 | 41.34 | 41.40 | 41.76 | 42.18 |
La Palma-VE-2021 | 61.59 | 62.89 | 66.47 | 65.58 | 62.75 | 65.67 | 65.36 |
Ukraine-AC-2022 | - | - | - | - | - | - | - |
Turkey-EQ-2023 | 48.91 | 48.87 | 50.93 | 50.93 | 53.12 | 52.60 | 54.14 |
Myanmar-CC-2023 | - | - | - | - | - | - | - |
Hawaii-WF-2023 | 46.68 | 54.29 | 47.71 | 58.36 | 60.40 | 60.25 | 59.43 |
Morocco-EQ-2023 | 42.41 | 40.54 | 42.62 | 42.98 | 43.15 | 43.91 | 44.43 |
Derna-FL-2023 | 50.05 | 53.17 | 60.50 | 59.40 | 58.45 | 62.86 | 64.68 |
Acapulco-HC-2023 | - | - | - | - | - | - | - |
Average | 49.71 | 50.79 | 50.98 | 54.14 | 54.08 | 56.18 | 55.45 |

### v2

@@ 3.3 Model training
To train these supervised models, we use a combination of cross-entropy loss and Lovasz softmax loss (Berman et al., 2018). Cross-entropy loss serves as the basic loss function for dense prediction tasks, while Lovasz softmax loss effectively addresses sample imbalance between non-building and building pixels and across different damage levels.
@@
For zero-shot and one-shot setups, we incorporate UDA and SSL techniques to better utilize limited or unlabeled target disaster data. The specifics of these training adjustments, along with the UMCD and UMIM approaches, are provided in Appendix E, F, and G, respectively.
| Beirut-EP-2020 | 52.23 | 48.53 | 51.85 | 52.04 | 52.57 | 60.55 | 49.38 |
| Bata-EP-2021 | 39.24 | 40.26 | 38.69 | 41.44 | 39.80 | 40.37 | 40.47 |
| Goma-VE-2021 | 61.52 | 59.26 | 62.95 | 62.50 | 61.84 | 61.55 | 62.89 |
| Les Cayes-EQ-2021 | 40.39 | 38.52 | 40.84 | 40.99 | 41.24 | 41.75 | 42.36 |
| La Palma-VE-2021 | 62.72 | 64.31 | 65.56 | 64.24 | 65.34 | 64.27 | 65.13 |
| Marshall-WF-2021 | 53.96 | 44.67 | 51.48 | 56.57 | 56.55 | 57.50 | 57.23 |
| Ukraine-AC-2022 | 43.92 | 43.84 | 40.71 | 45.52 | 45.56 | 47.45 | 47.43 |
| Turkey-EQ-2023 | 49.34 | 49.53 | 48.57 | 50.69 | 52.34 | 51.36 | 52.70 |
| Kyaukpyu-CC-2023 | 38.11 | 38.56 | 40.49 | 42.35 | 42.03 | 48.42 | 41.05 |
| Hawaii-WF-2023 | 53.02 | 57.81 | 57.92 | 55.34 | 58.40 | 60.82 | 60.70 |
| Morocco-EQ-2023 | 42.62 | 42.05 | 42.58 | 43.17 | 43.05 | 44.00 | 44.06 |
| Derna-FL-2023 | 59.47 | 56.08 | 66.82 | 59.48 | 57.73 | 62.24 | 63.30 |
| Acapulco-HC-2023 | 49.21 | 46.66 | 42.15 | 49.98 | 51.08 | 50.82 | 47.01 |
| Noto-EQ-2024 | 36.49 | 37.33 | 36.56 | 38.35 | 42.84 | 40.57 | 45.78 |
| Average | 48.73 | 47.67 | 49.08 | 50.19 | 50.74 | 52.26 | 51.39 |
@@ 3.4 Accuracy assessment
We adopt overall accuracy (OA), F1 score (F1), and mean intersection over union (mIoU) to evaluate the performance of DL models on building damage assessment. These metrics are commonly used in previous research on building damage assessment (Zheng et al., 2021). Following the setup in prior unimodal building damage assessment studies and the related xView2 Challenge (Gupta et al., 2019), F1 score is used to assess the performance of DL models in the building localization and damage classification subtasks. Finally, OA and mIoU are employed to measure the overall quality of the proxy damage map, providing a comprehensive assessment of the models’ ability to localize buildings and classify damage levels accurately.
4 Results and Analysis
4.1 Evaluation results on standard machine learning split
Table 5 shows the evaluation results for each model on the test set. We observe that ChangeMamba achieves the best overall performance, with an OA of 96.22, an mIoU of 67.63, and the highest and scores (90.90 and 72.70, respectively). DamageFormer also performs well, closely following ChangeMamba, with an mIoU of 67.09 and an OA of 96.13. Both models demonstrate strong capability in both building localization and damage classification tasks. The accuracy advantage of ChangeMamba and DamageFormer underscores the importance of leveraging advanced deep learning architectures to improve performance in complex tasks such as building damage assessment. For models using the direct prediction approach (UNet, DeepLabV3+, SiamAttnUNet), UNet delivers the best results, with an mIoU of 64.94 and an OA of 95.47. However, its performance still lags behind that of the decoupled models, further emphasizing the advantage of task decoupling.
To ensure that the evaluation is not dominated by a few events with a large number of images, e.g., Turkey-EQ-2023, Table 6 further presents the event-level mIoU for each model across 14 disaster events. ChangeMamba and DamageFormer achieve the highest average mIoU, with scores of 51.39 and 52.26, respectively. Notably, DamageFormer performs particularly well on events such as Beirut-EP-2020, Marshall-WF-2021, and Derna-FL-2023, demonstrating its robustness across different types of disasters. While performance varies across events, earthquake-related events such as Les Cayes-EQ-2021, Morocco-EQ-2023, and Noto-EQ-2024 present greater challenges for all models, where the average mIoU remains relatively low. This highlights the need for further research to improve model robustness for earthquake damage assessment, where damage patterns are often more complex and diverse.
@@ 4.1 Evaluation results on standard machine learning split
4.2 What have models learned and what can they learn?
To better understand model behavior beyond performance metrics, we explore the internal attention patterns of trained models using class activation maps (CAMs). Figure 7 presents the CAM responses of ChangeOS across six representative disaster events, including volcano, explosion, flood, wildfire, earthquake, and hurricane. We observe that the attention distribution varies across disaster types. In Goma-VE-2021, for the “Destroyed” category, the encoder processing the optical imagery exhibits strong activations across nearly all built-up regions, accurately localizing individual buildings. This suggests the model has effectively learned to extract detailed structural cues from pre-event optical imagery. In contrast, the encoder handling the SAR imagery shows intense activation over the lava-covered region on the left, indicating that the model has identified this as a key signal for destruction—likely due to the significant backscatter changes caused by lava flow. In the “Damaged” category, however, the activations are more subtle. Attention is primarily focused near the boundary of the lava flow, where partial or ambiguous structural changes occur. In the decoder stage, the destroyed buildings in the lava-affected area are strongly activated, aligning well with the reference labels. Conversely, for the “Damaged” class, only a few regions are activated. This suggests that inferring partial damage from SAR imagery in volcanic disaster scenarios remains a significant challenge, as the subtle structural degradation is often not clearly reflected in SAR backscatter or texture.
Beyond understanding what the models have already learned, a more critical question is: What can they learn? More specifically, to what extent can optical–SAR combinations support accurate building damage assessment across diverse disaster scenarios? To investigate this, we aggregate the mIoU of the seven models and show their average mIoU across seven major disaster types in Figure 8. Most notably, wildfire and volcano events exhibit the highest IoU scores for the “Destroyed” category, both exceeding 70. This indicates that the model can effectively leverage SAR-based backscatter anomalies, such as lava flows or widespread debris fields, to detect fully destroyed structures. These results demonstrate the strong potential of Optical–SAR fusion in such high-impact scenarios. However, for the “Damaged” category, the performance drops significantly. The average IoU for damaged buildings in wildfire events remains below 20, while in volcano events, it falls to around 5. This suggests that single-polarization SAR imagery lacks the fine-grained information needed to reliably distinguish partially damaged buildings, where structural integrity may still be partially preserved and backscatter signals remain ambiguous. For earthquake events, both damaged and destroyed categories yield relatively low IoUs. This is likely due to the complex and heterogeneous patterns of structural collapse typical of seismic events, where damage is often subtle, partial, and highly variable. These conditions pose significant challenges for SAR-based assessment. Interestingly, for flood and hurricane events, the model achieves relatively high IoU scores for the “Damaged” category with about 50 and 60, highlighting that SAR is effective at capturing contextual environmental changes—such as water inundation or terrain disruption—which indirectly aid in assessing building damage. In the case of the conflict event, the model’s performance on the destroyed class is surprisingly low. We attribute this to the limited number of destroyed samples in the dataset for this category, which leads to insufficient learning and poor generalization.
In summary, these findings confirm both the promise and limitations of Optical–SAR combinations for all-weather, global-scale disaster response. While the combination performs well in events characterized by large-scale surface disruption (e.g., wildfires, volcanoes), it struggles with subtle or localized damage patterns. Incorporating richer data sources, such as fully polarimetric SAR and LiDAR, can further enhance the accuracy and reliability of future building damage assessments.
| Method | mIoU (set) | mIoU (event) |
| Baseline | 67.63 | 51.39 |
| Test-time augmentation | 68.50 0.87 | 51.95 0.56 |
| Object-based major voting | 67.22 0.41 | 52.08 0.69 |
| Ensembling multiple models | 68.45 0.82 | 52.14 0.75 |
| All | 68.86 1.23 | 52.31 0.92 |
4.3 Effect of post-processing methods
Post-processing techniques help refine the raw predictions from deep learning models, reducing noise, improving consistency, and ensuring spatial coherence in damage proxy maps (Zheng et al., 2021). Here, we further explore the impact of post-processing algorithms. Table 7 presents the effect of three post-processing techniques applied to ChangeMamba. The post-processing methods evaluated include test-time augmentation, object-based major voting, and model ensembling. Further details on these methods are provided in Appendix D
Firstly, test-time augmentation improves the mIoU by 0.87 at the set level and 0.56 at the event level, demonstrating its effectiveness in enhancing model robustness across diverse disaster scenarios. Object-based major voting, which aggregates predictions at the building-object level to enforce spatial consistency, slightly reduces set-level mIoU (-0.41) but improves event-level mIoU (+0.69). Ensembling multiple models leads to a 0.82 increase in mIoU at the set level and a 0.75 increase at the event level, reinforcing its effectiveness in improving model performance across different disaster events. Applying all post-processing techniques together yields the highest performance improvement, with a 1.23 increase in set-level mIoU and a 0.92 increase in event-level mIoU. These results confirm that a combination of different post-processing methods can significantly enhance the reliability of AI-driven damage assessments, ensuring better generalization across disaster types and locations.
In summary, post-processing techniques play a crucial role in improving damage proxy maps. Future work can further explore adaptive post-processing strategies tailored to specific disaster types to enhance prediction reliability in multimodal EO data contexts.
4.4 Evaluation results on cross-event transfer split
4.4.1 Baseline methods
@@
| zero- / one-shot | zero- / one-shot | zero- / one-shot | zero- / one-shot | zero- / one-shot | zero- / one-shot | zero- / one-shot | |
| Beirut-EP-2020 | 31.62 / 34.89 | 34.72 / 33.58 | 35.33 / 29.17 | 39.43 / 40.94 | 38.48 / 38.21 | 38.71 / 39.93 | 40.30 / 40.92 |
| Bata-EP-2021 | 33.48 / 37.48 | 33.55 / 30.23 | 33.78 / 35.24 | 37.46 / 40.06 | 38.00 / 38.18 | 38.91 / 39.42 | 39.17 / 40.31 |
| Goma-VE-2021 | 55.97 / 57.14 | 55.01 / 58.13 | 57.52 / 55.75 | 58.91 / 59.54 | 58.66 / 57.85 | 55.86 / 59.89 | 60.39 / 61.91 |
| Les Cayes-EQ-2021 | 26.21 / 38.72 | 35.50 / 37.90 | 37.22 / 38.76 | 39.45 / 40.28 | 39.36 / 40.26 | 40.41 / 41.07 | 41.35 / 42.06 |
| La Palma-VE-2021 | 32.74 / 35.29 | 30.86 / 36.25 | 34.00 / 34.99 | 32.90 / 33.32 | 29.96 / 32.60 | 32.28 / 37.62 | 34.18 / 34.93 |
| Marshall-WF-2021 | 34.97 / 37.84 | 35.05 / 37.12 | 35.23 / 39.54 | 39.82 / 46.83 | 36.31 / 43.72 | 40.93 / 51.15 | 40.61 / 45.96 |
| Ukraine-AC-2022 | 32.36 / 37.50 | 32.23 / 35.98 | 33.84 / 35.76 | 34.77 / 38.14 | 36.30 / 39.28 | 33.84 / 36.51 | 35.78 / 36.99 |
| Turkey-EQ-2023 | 37.82 / 35.83 | 37.28 / 40.49 | 36.09 / 37.49 | 38.32 / 40.11 | 38.45 / 39.51 | 37.20 / 40.27 | 41.28 / 42.80 |
| Kyaukpyu-CC-2023 | 31.46 / 31.71 | 34.01 / 34.75 | 33.77 / 35.93 | 34.19 / 37.71 | 36.66 / 36.25 | 32.99 / 33.54 | 33.85 / 36.08 |
| Hawaii-WF-2023 | 31.73 / 39.80 | 31.32 / 45.44 | 31.79 / 45.12 | 30.94 / 42.65 | 31.15 / 41.96 | 31.60 / 50.64 | 36.14 / 50.89 |
| Morocco-EQ-2023 | 38.24 / 39.81 | 40.48 / 41.23 | 40.78 / 40.69 | 40.89 / 42.89 | 41.40 / 41.89 | 41.89 / 43.42 | 42.86 / 43.51 |
| Derna-FL-2023 | 33.25 / 38.44 | 35.07 / 40.02 | 33.76 / 42.15 | 35.88 / 44.77 | 35.80 / 43.85 | 36.27 / 44.69 | 37.02 / 45.47 |
| Acapulco-HC-2023 | 26.99 / 38.76 | 27.78 / 33.24 | 30.93 / 38.94 | 26.70 / 34.53 | 27.06 / 34.50 | 27.48 / 32.87 | 28.45 / 38.96 |
| Noto-EQ-2024 | 31.02 / 33.83 | 35.57 / 39.17 | 35.87 / 36.22 | 36.04 / 38.52 | 39.14 / 39.15 | 36.37 / 43.93 | 39.90 / 44.33 |
| Average | 34.13 / 38.36 | 35.60 / 38.82 | 36.42 / 38.98 | 37.55 / 41.45 | 37.62 / 40.52 | 37.48 / 42.49 | 39.38 / 43.23 |
Cross-event transfer, especially under zero-shot settings, poses a significant challenge for building damage assessment. As shown in Table 8, the average mIoU across all baseline models in the zero-shot setting is below 40, a noticeable drop compared to their performance under the fully supervised standard ML split, where models typically achieve 48 to 52 mIoU. This performance gap underscores the difficulty of generalizing to unseen disaster events without access to any target-domain supervision due to substantial domain shifts in imaging conditions, damage patterns, urban morphology, and sensor response. Despite this, all models exhibit clear performance gains in the one-shot setting, where a small number of labeled samples from the target event are available. This suggests that even minimal supervision can significantly aid adaptation to new disaster contexts.
Among the evaluated models, ChangeMamba consistently achieves the highest overall performance, with an average mIoU of 39.38 in zero-shot and 43.23 in one-shot settings. It is closely followed by DamageFormer, highlighting the strength of recent transformer- and Mamba-based architectures in transferring learned knowledge under complex multimodal and disaster scenarios. Models using a decoupled architecture generally outperform the direct segmentation models, confirming that separating building localization and damage classification improves generalization in cross-event transfer tasks. Conversely, events like Acapulco-HC-2023 and Ukraine-AC-2022 show weaker results across all models, reflecting the difficulty in transferring to domains with limited or inconsistent destruction patterns.
4.4.2 Unsupervised domain adaptation and semi-supervised learning methods
| Method | Avg. mIoU |
| Source-Only | 35.60 |
| AdaptSeg (Tsai et al., 2018) | 36.05 0.45 |
| AdvEnt (Vu et al., 2019) | 35.43 0.17 |
| CLAN (Luo et al., 2019) | 36.01 0.41 |
| PyCDA (Lian et al., 2019) | 33.19 2.41 |
| FDA (Yang and Soatto, 2020) | 36.29 0.69 |
| Method | Avg. mIoU |
| Baseline | 38.82 |
| MT (Tarvainen and Valpola, 2017) | 40.00 1.18 |
| CCT (Ouali et al., 2020) | 39.49 0.67 |
| GCT (Ke et al., 2020) | 39.34 0.52 |
| CPS (Chen et al., 2021) | 37.06 1.76 |
In the context of cross-event transfer, UDA and SSL methods naturally emerge as promising strategies to bridge the domain gap between source and target disaster events. In the zero-shot setting, UDA methods aim to improve model generalization by aligning the source and target domains without requiring any target labels. In the one-shot setting, SSL methods further leverage a small number of labeled samples along with abundant unlabeled data from the target event, making them especially appealing for real-world disaster scenarios where rapid annotation is infeasible. However, it remains an open question whether these UDA and SSL methods, originally developed for natural image domains, can effectively handle the challenges of multimodal EO data in complex disaster scenarios. To examine the applicability of these methods, we evaluate several representative UDA and SSL algorithms using DeepLabV3+ as the baseline model. The results are reported in Tables 9 and 10, respectively.
Table 9 reports the performance of five classical UDA methods on the BRIGHT dataset. Compared to the source-only baseline (35.60 mIoU), most methods achieve modest improvements, with FDA (+0.69) and AdaptSeg (+0.45) showing the most consistent gains. CLAN also performs slightly better (+0.41), suggesting that category-level alignment contributes positively even under a large domain shift. In contrast, AdvEnt shows negligible change, and PyCDA significantly underperforms, dropping over 2 mIoU compared to the baseline. This performance degradation indicates that approaches relying on pseudo-label refinement or curriculum learning may struggle in disaster imagery, where spatial layout and damage semantics vary drastically across events. Overall, while UDA methods show some promise, their improvements are relatively minor and not robust across all events, underscoring the difficulty of domain alignment in EO-based damage mapping.
In the one-shot setting, we evaluate several popular SSL methods, summarized in Table 10. With only minimal supervision in the target event, all methods except CPS achieve consistent improvements over the supervised one-shot baseline (38.82). The best-performing method is Mean Teacher (MT), which yields a gain of +1.18, followed by CCT (+0.67) and GCT (+0.52). These results show that simple consistency-based teacher-student frameworks are particularly effective in leveraging unlabeled data under limited supervision, likely due to their robustness to noisy or class-imbalanced targets.
4.4.3 Why cross-event transfer is challenging?
To better understand why cross-event generalization is so difficult, we explore two fundamental factors rooted in the nature of real-world disaster response.
1.
Inconsistent damage signatures across events. Figure 9 presents violin plots of SAR backscatter values for intact, damaged, and destroyed buildings across multiple event pairs. These plots reveal two key observations. Firstly, even within the same disaster category (the first row), the pixel intensity distributions for damaged and destroyed buildings differ significantly between events. This indicates that SAR-based damage signatures are not consistent across locations, possibly due to differences in urban layout, building materials, or sensor incidence angles. In addition, the distribution shift becomes more pronounced across different disaster types (the second row). For instance, the destroyed-building signatures in wildfires are notably different from those in volcano eruptions, which again differ from floods or hurricanes. These variations reflect fundamental differences in damage mechanisms: buildings burned in wildfires, submerged in floods, or collapsed in earthquakes leave very different patterns in SAR backscatter. Such distributional discrepancies make it extremely difficult for models to generalize from one event to another. A model trained on one disaster may learn class boundaries (e.g., between “Damaged” and “Destroyed”) that do not transfer well to another disaster, especially when the visual and physical properties of damage are fundamentally different.
Figure 10: Comparison of models’ best performance (mIoU) on test events versus the best checkpoints selected on validation sets under the cross-event transfer setting. The farther a point lies from the diagonal line, the larger the gap between the model’s selected performance and its true upper bound. -
2.
Lack of target domain supervision for model selection. Another critical challenge in cross-event setup is the absence of labeled target-domain samples for model selection. In typical domain adaptation benchmarks in computer vision, a target domain validation set is available to tune hyperparameters or select the best checkpoint. However, in disaster response scenarios, no such validation data can be assumed. Figure 10 illustrates the resulting issue: the gap between the actual best mIoU on the test event and the mIoU of the checkpoint selected using a source-domain validation set for several representative models under both zero-shot and one-shot settings. In the zero-shot setting, this performance gap is substantial across most models. This indicates that relying on source-domain validation leads to suboptimal model selection due to poor reflection of the target distribution. The one-shot setting helps reduce this gap by enabling limited target-aware selection. However, the issue is not entirely resolved. Even with a few labeled samples, model instability and domain shift still make selection challenging.
Together, these two findings highlight that cross-event transfer is hindered not only by semantic and statistical shifts across disaster types but also by operational constraints that prevent ideal training and tuning. Future solutions should consider both aspects—developing models that are robust to distributional variance and designing selection strategies that do not depend on target supervision, such as self-validation, early stopping heuristics, or domain-agnostic performance proxies (Yang et al., 2023).
4.5 Extended evaluation of unsupervised multimodal change detection methods
| Method | UMCD benchmark datasets | BRIGHT | |||||
| OA | F1 | KC | OA | F1 | IoU | KC | |
| IRG-McS (Sun et al., 2021) | 98.3 / - / 97.1 / 97.2 | 80.4 / - / 75.4 / 73.7 | 79.4 / - / 73.9 / 75.1 | 90.03 | 12.65 | 6.75 | 7.65 |
| SR-GCAE (Chen et al., 2022b) | 98.6 / 98.5 / - / - | 82.9 / 77.6 / - / - | 82.1 / 76.9 / - / - | 77.83 | 14.35 | 7.73 | 5.64 |
| FD-MCD (Chen et al., 2023) | 98.2 / 97.8 / - / 96.7 | 81.4 / 72.2 / - / 73.2 | 82.3 / 71.1 / - / 71.4 | 80.96 | 15.84 | 8.60 | 7.94 |
| AOSG (Han et al., 2024) | - / - / - / 96.4 | - / - / - / 77.7 | - / - / - / 75.9 | 77.93 | 10.75 | 5.68 | 3.98 |
| AGSCC (Sun et al., 2024a) | 98.3 / - / 95.9 / 97.6 | 78.2 / - / 68.0 / 77.9 | 77.3 / - / 65.8 / 76.6 | 88.49 | 14.82 | 8.00 | 9.54 |
| AEKAN (Liu et al., 2025) | 98.7 / - / - / 98.3 | 83.8 / - / - / 84.7 | 83.1/ - / - / 83.9 | 81.60 | 13.09 | 7.00 | 3.56 |
Unsupervised multimodal change detection (UMCD) plays a crucial role in post-disaster assessment by enabling rapid analysis of affected areas without requiring labeled data or prior model training (Lv et al., 2022). The current research is mainly conducted on toy datasets with limited geographic diversity and scene complexity. These datasets mainly involve general land-cover changes. It is still unknown whether these methods can work in real disaster occurrence/building damage scenarios. BRIGHT offers a new opportunity to evaluate UMCD methods, thereby providing insights into their robustness and scalability. Six representative UMCD methods (Sun et al., 2021; Chen et al., 2022b, 2023; Han et al., 2024; Sun et al., 2024a; Liu et al., 2025) are evaluated in this paper. The specific experimental setup is described in Appendix G.
Table 11 reports the performance of these methods on BRIGHT and their accuracy on current UMCD benchmark datasets shown in Figure 12. Although these methods can achieve considerable performance on the current benchmark dataset, their performance suffers noticeable declines on BRIGHT. For instance, while these methods achieve F1 scores between 70-85 on existing benchmarks, their F1 scores drop below 20 on BRIGHT. This dramatic performance gap underscores the limitations of current UMCD research and highlights the challenges posed by real disaster scenarios. We identify three primary reasons for this decline:
1.
Limitations in traditional UMCD datasets: The traditional UMCD datasets consist of only a handful of image pairs, often depicting simple land-cover changes such as urban expansion or deforestation, with relatively low-resolution imagery with limited geographic diversity, as shown in Figure 12. These datasets fail to capture the complexity and variability found in real-world disaster scenarios. In contrast, BRIGHT provides thousands of high-resolution multimodal image pairs across different types of disasters, significantly increasing the diversity of test cases and making it a more challenging benchmark for UMCD model evaluation.
2.
Interference from non-disaster changes: Unlike previous UMCD benchmarks where general land-cover changes are the primary detection target, changes such as vegetation growth, water body shifts, and urban development may introduce noise and interfere with damage detection in real disaster scenarios. While prior UMCD studies treat such changes as valid detection targets, BRIGHT requires methods to differentiate true building structural damage from irrelevant land-cover variations, posing a unique challenge to current approaches.
3.
Problematic evaluation protocol: Current UMCD research follows a problematic evaluation protocol where models are trained, hyperparameters are tuned, and then validated on the same dataset. This approach can actually lead to overfitting, artificially making the model look very accurate but not generalizable to real-world scenarios. In a real disaster scenario, there are no labeled samples available for hyperparameter tuning, making the current practice unrealistic. BRIGHT exposes this limitation, as models now need to be trained or tuned before being tested directly on new, unseen data, requiring truly generalizable and adaptive learning strategies.
The results highlight the challenges of applying existing UMCD models to real-world disaster scenarios. The BRIGHT dataset reveals significant limitations in current methodologies and also presents new opportunities for future UMCD research.
4.6 Extended evaluation of unsupervised multimodal image matching methods
| Method | Before (meters) | After (meters) |
| FLSS (Ye et al., 2017) | 16.97 | 10.01 |
| HOPC (Ye et al., 2019) | 9.37 | |
| LNIFT (Li et al., 2022) | N/A | |
| SRIF (Li et al., 2023a) | N/A |
Finally, precise image alignment is a critical prerequisite for any multimodal EO application. BRIGHT offers a unique opportunity to evaluate the performance of existing UMIM algorithms under realistic, large-scale disaster conditions. Due to the lack of real pixel-level ground truth correspondences, we adopt a proxy evaluation strategy using manually selected control points (as shown in Figure 11) as references. These points are selected by several EO experts to represent clearly identifiable and stable features across modalities. While this does not constitute absolute ground truth, the use of such human-verified correspondences provides a valuable reference, allowing us to assess how closely automated methods approximate human matching ability under multimodal and disaster conditions.
Table 12 presents the quantitative results of several representative UMIM methods on the Noto-Earthquake-2024 scene. Our experimental results reveal that traditional feature-based methods generally fail to achieve successful automatic registration. We attribute this limitation to the large spatial extent of the scenes and the drastic cross-modal differences, which make direct keypoint matching highly error-prone and unreliable. In contrast, area-based methods, which rely on region-level similarity rather than point-level correspondence, can partially mitigate registration errors and achieve relatively more stable performance. However, these methods still face a critical limitation: they typically perform local matching within a fixed search window. This strategy is inherently inadequate for handling large displacements, especially in SAR imagery over mountainous or uneven terrain, where terrain-induced distortions can cause substantial pixel shifts that far exceed the local search range. Thus, automatic methods may serve as a useful auxiliary tool for preliminary matching or candidate generation, but human expertise remains indispensable for ensuring precise and reliable alignment in operational scenarios. We hope that the BRIGHT benchmark and its challenging real-world settings will inspire the development of new UMIM methods that are more robust to large scene variations, sensitive to terrain-induced distortions, and ultimately capable of reducing human efforts in future operational EO-based disaster response workflows.
5 Discussion